# MSDNet: Multi-Scale Decoder for Few-Shot Semantic Segmentation via Transformer-Guided Prototyping

### v4

@@ Abstract
Keywords:
Few-shot learning, few-shot segmentation, Semantic Segmentation, Prototype generation
@@ 3.2.2 Support Prototype
Where $F_{pool}$ represents the average pooling operation, and $\odot$ signifies element-wise multiplication (Hadamard product) with the support mask $M_{s}$. The MAP operation involves computing the average pooling of the masked feature map, focusing solely on regions of interest specified by the support mask. This results in the generation of the Support Prototype, which encapsulates essential semantic information from the support example, facilitating effective few-shot segmentation.
@@ 3.2.4 Multi Scale Decoder
Notably, one of the distinctive aspects of our multi-scale decoder is the incorporation of mid-level and high-level features from the encoder, like U-Net architecture. Specifically, in each stage of the decoder, the input to the residual layers combines the output from the previous layer with either the $conv5\_x$ features (the output of the last block of the encoder) or the merged mid-level features ($X_{s}^{merged}$) extracted from the support image’s encoder. This fusion of features from different levels of abstraction enhances the decoder’s ability to capture both detailed and contextual information essential for accurate segmentation.
@@ 3.3 Loss function
$Dice\;Loss=1-\frac{2\times\left|M\bigcap M_{q}\right|}{\left|M\right|+\left|M_{q}\right|}$ (4)
Where $\left|M\bigcap M_{q}\right|$ represents the intersection between the predicted and ground truth masks, and $\left|M\right|$ and $\left|M_{q}\right|$ denote the cardinality of the predicted and ground truth masks, respectively. Minimizing the Dice loss encourages the model to generate segmentation masks that closely match the ground truth masks, leading to more accurate segmentation results during training.
@@ 4.1 Datasets
[…]
@@ 4.2 Experimental Setting
Backbone Methods Publication fold0 fold1 fold2 fold3 mean FB-IoU fold0 fold1 fold2 fold3 mean FB-IoU # learnable params
@@
Backbone Methods Publication 1-shot 5-shot # learnable params
@@ 4.4 Comparison with SOTA
[…]
@@
Table 3: Few-shot segmentation performance on cross-dataset task, ”$COCO-20^{i}\rightarrow PASCAL-5^{i}$”, in terms of mIoU, with different backbones (ResNet-50 and ResNet-101). Numbers in bold represent the best performance, while underlined values denote the second-best performance.
@@ 4.6 Ablation Study
Baseline CMGM STD Multi Scale Decoder 1-shot
@@
# residual blocks 1-shot # learnable params

### v5

@@ 3.2.2 Support Prototype
Where $F_{pool}$ represents the average pooling operation, and signifies element-wise multiplication (Hadamard product) with the support mask $M_{s}$. The MAP operation involves computing the average pooling of the masked feature map, focusing solely on regions of interest specified by the support mask. This results in the generation of the Support Prototype, which encapsulates essential semantic information from the support example, facilitating effective few-shot segmentation.
@@ 3.2.4 Multi Scale Decoder
Notably, one of the distinctive aspects of our multi-scale decoder is the incorporation of mid-level and high-level features from the encoder, like U-Net architecture. Specifically, in each stage of the decoder, the input to the residual layers combines the output from the previous layer with either the conv5_xconv5\_x features (the output of the last block of the encoder) or the merged mid-level features ($X_{s}^{merged}$) extracted from the support image’s encoder. This fusion of features from different levels of abstraction enhances the decoder’s ability to capture both detailed and contextual information essential for accurate segmentation.
@@ 3.3 Loss function
[…]
@@ 4.1 Datasets
$\textbf{PASCAL-5}^{\textbf{i}}$ Dataset. The $PASCAL-5^{i}$ dataset, introduced by Shaban et al. shaban2017one, is derived from the PASCAL VOC dataset everingham2010pascal, and augmented with the SDS hariharan2011semantic. The original PASCAL VOC dataset comprises 20 object categories. For $PASCAL-5^{i}$, these 20 categories are evenly divided into 4 subsets, each denoted as $PASCAL-5^{i}$. Consequently, each subset consists of 5 distinct object categories.
$\textbf{COCO-20}^{\textbf{i}}$ Dataset. The $COCO-20^{i}$ dataset, introduced by Nguyen et al. nguyen2019feature, is derived from MSCOCO dataset lin2014microsoft. The $COCO-20^{i}$ dataset includes a total of 80 object categories. Similar to $PASCAL-5^{i}$, these 80 categories are divided into 4 subsets, with each subset denoted as $COCO-20^{i}$. Each subset contains 20 distinct object categories. Notably, $COCO-20^{i}$ presents a greater challenge due to its larger number of categories and images compared to $PASCAL-5^{i}$.
@@ 4.2 Experimental Setting
Backbone Methods Publication fold0 fold1 fold2 fold3 mean FB-IoU fold0 fold1 fold2 fold3 mean FB-IoU
# learnable
params
@@
Backbone Methods Publication 1-shot 5-shot
# learnable
params
@@ 4.4 Comparison with SOTA
Results on $\textbf{PASCAL-5}^{\textbf{i}}$ Dataset. As shown in Table 1, our proposed method, utilizing ResNet50 and ResNet101 backbones, consistently surpasses SOTA methods in both 1-shot and 5-shot scenarios across all four folds of the $PASCAL-5^{i}$ dataset. Notably, our method achieves competitive performance across all folds, frequently ranking among the top-performing methods in both 1-shot and 5-shot settings.
Results on $\textbf{COCO-20}^{\textbf{i}}$ Dataset. Similarly, Table 2 presents the results on the $COCO-20^{i}$ dataset, where our proposed method demonstrates strong performance under both ResNet50 and ResNet101 backbones across 1-shot and 5-shot settings. In many folds, our approach achieves the highest or second-highest mIoU scores, reflecting its robustness and efficiency. Additionally, our model achieves competitive mean and FB-IoU scores while maintaining a significantly smaller number of learnable parameters compared to other methods.
@@
Table 3: Few-shot segmentation performance on cross-dataset task, $COCO-20^{i}\rightarrow PASCAL-5^{i}$”, in terms of mIoU, with different backbones (ResNet-50 and ResNet-101). Numbers in bold represent the best performance, while underlined values denote the second-best performance.
@@ 4.6 Ablation Study
Baseline CMGM STD
Multi Scale
Decoder
1-shot
@@
# residual
blocks
1-shot
# learnable
params